PROBLEM SOLVING INVOLVING THE NORMAL CURVE CONCEPTS


Given = 30 and = 4.5. Express the raw score for each of the following.

1. z = 1.25

2. z = -1.67

3. z = 2.3

4. z = -0.30

5. z = 1.96


1
Expert's answer
2021-05-07T10:38:57-0400

Given, "\mu=30,\sigma=4.5"


1. z = 1.25


Raw score "x=\mu+z\times \sigma=30+1.25\times 4.5=35.625"


2. z = -1.67


Raw score "x=\mu+z\times \sigma=30-1.67\times 4.5=22.485"

3. z = 2.3


Raw score "x=\mu+z\times \sigma=30+2.3\times 4.5=40.35"


4. z = -0.30


Raw score "x=\mu+z\times \sigma=30-0.30\times 4.5=28.65"


5. z = 1.96


Raw score "x=\mu+z\times \sigma=30+1.96\times 4.5=38.82"